Can you please help me with this??
root 3x+4 - root 2x-4=2

Answer by oscargut(2103)

root 3x+4 - root 2x-4=2
taking squares
(root 3x+4 - root 2x-4)^2=4
using(a+b)^2=a^2+2ab+b^2
3x+4+2root(3x+4)root(2x-4)+2x-4=4
5x-4=2root(3x+4)root(2x-4)
(5x-4)^2=4(3x+4)(2x-4)
25x^2-40x+16=24x^2-16x-64
x^2-24x+80=0
so solutions are 4 and 20 (check it at the original equation)
Answer: solutons are x=4 and x=20